Х Evaluate dx x? + 1
use the fundalmental therom of Calculus if applicable
Answer #1

To evaluate \int \frac{x}{\sqrt{x^2+1}}dx make the substitution x=\tan t then dx=\sec^2 tdt .

\int \frac{x}{\sqrt{x^2+1}}dx=\int \frac{\tan t}{\sqrt{\tan^2 t+1}}\sec^2 tdt\\ \int \frac{x}{\sqrt{x^2+1}}dx=\int \frac{\tan t}{\sec t}\sec^2 tdt\\ \int \frac{x}{\sqrt{x^2+1}}dx=\int \tan t\sec tdt\\ \int \frac{x}{\sqrt{x^2+1}}dx=\sec t+C\\ {\color{Blue} \int \frac{x}{\sqrt{x^2+1}}dx=\sqrt{x^2+1}+C}